lavnotlez:

In a society roads and buildings are created. Find the total area covered by roads and buildings. Find the cost of building roads if it costs 3000 per m² ​

lavnotlez:

1 attachment
Florisalreadytaken:

I am not sure what currency that is so we will be using $ okay. we are given this rectangle which supposedly represents an area made of 5 same roads and 4 other rectangles which one would assume they're equal but we don't have any information about that. It's asking you to find the total area so that would be: \( A_t = 30 m \times 50 m \) \( A_t = (30 \times 50)m^2 \) \( \boxed{A_t = 1500 \ m^2 }\) The second question wants you to find the cost of building roads which I am not sure whether they mean one road or all \(5\) of them? this is the solution for one road: \( A_{R} =3 m \times 30 m \) \( A_{R} = (3 \times 30)m^2 \) \( \boxed{A_{R} = 90 m^2 }\)


this is the solution for \(5\) roads: \( A_{5R} =5\left(3 m \times 30 m \right) \) \( A_{5R} =5\left[ (3 \times 30)m^2 \right] \) \( A_{5R} =5 \times 90m^2 \) \( \boxed{A_{5R} = 450 m^2 }\)

Florisalreadytaken:

Now, can you find the price having the areas?